# Time-dependent conserved operators for autonomous systems and quantization of resistance

### v2

@@ Abstract
Two systems for a charged particle are studied, the first one when it is under the effect of a constant electric field, and the second one when it is under the effect of a constant electromagnetic field. For both systems, it is possible to find time-dependent conserved operators that can be used to derive time-dependent wave functions to the complete Schrödinger equation, such that the time variable is not separable from the space coordinates. At the same time, these conserved operators are used to build up a unitary operators, which define the symmetries of the systems. Then, it is shown that the invariance of the wave function under the action of these unitary operators leads to the quantization of resistance as integer multiples of the Klitzing’s constant.
Time-dependent operators, Resistance quantization, Klitzing’s constant
@@ I Introduction
Conserved operators in dynamical systems have proven to be valuable tools in quantum mechanics for identifying system bases. For example, in the central force problem, angular momentum is conserved and used to determine the eigenfunctions associated with azimuthal and polar angles [1]. Another illustrative case involves a constant magnetic field, where, using Landau’s gauge, linear momentum is conserved and used to find the system’s basis. This is known as Landau’s ansatz, and the resulting basis turns out to be the product of plane waves and the harmonic oscillator wave function [2]. Recently, for a constant electromagnetic field, it has been shown that, in addition to Landau’s ansatz, a time-dependent conserved operator is necessary for a complete description of the system [3]. In the latter work mentioned, it was demonstrated that the invariance of the wave function under the action of a unitary transformation leads to the quantization of resistivity for a single particle, yielding a result similar to the experimentally measured expression by Störmer [4, 5, 6]. Furthermore, this invariance condition leads to two quantized quantities: the magnetic flux quantization (also derived using Landau’s ansatz [2]) and a second quantization condition related to the electric field, given by Eq. (34). The latter condition is the principal point of this work.
@@
Two autonomous systems exhibiting time-dependent conserved operators are presented here. The first system involves a particle under the influence of a constant electric field, while the second describes a particle subjected to a constant electromagnetic field. The system with a linear potential is of special interest when studying the shift in energy levels (Stark effect) [7, 8, 9, 10] and time-dependent solutions have been found using the Baker-Campbell-Hausdorff generalized formula for arbitrary initial conditions [11]. However, the approach presented in this paper differs from the previous one by being based on the conserved properties of the system, also known as the Lewis and Riesenfeld invariant method. Although it was initially developed to address non-autonomous systems [12, 13], here it is shown that it can also be applied to autonomous systems. On the other hand, the second system when a constant magnetic field is added such that is parallel to the electric field had been used to study the photodetachment from negative ions and the spatial distribution of monoenergetic electrons [14, 15]. However, aside from these applications, the main result of this work is to demonstrate that these time-dependent wave functions require the quantization of resistance, which is proportional to an integer multiple of Klitzing’s constant [16, 17, 18], when they are invariant under the action of a unitary transform defined by the time-dependent conserved operator.
II About time-dependent conserved operators
The Hamiltonian we are interested in is written in CGS units as
@@ III One-dimensional constant electric field
where, for simplicity, the eigenvalues were written as $-q{\cal E}\delta t$ such that $\delta t$ is a constant. Solving the above equation, it can be found that
@@
where the normalization was calculated by integrating in a length of $x\in[-L/2,L/2]$. At this point, one can search for the simplest form of the solution, in a sense that the above wave function is just a unitary transform of the new function. The conserved operators in Eq.(10) and Eq.(11) can be used to define a couple of unitary operators
@@
where $\delta x$ is a constant. These operators define the symmetries of the system such that the Schrödinger equation remains invariant under a unitary transformation, i.e.,
@@ IV Constant electromagnetic field
As mentioned in the introduction, another autonomous system with time-dependent conserved operators is the case of a constant electromagnetic field. The analysis for the situation where the electric field is perpendicular to the magnetic field can be found in reference [3]. Here, it will be consider the scenario where the electric field is parallel to the magnetic field. In this case, the same electric potential than in the previous section is used. On the other hand, to consider the magnetic field in Hamiltonian Eq.(1), it is necessary to define the vector potential. In this case ${\bf A}=B(0,0,y)$ satisfies the requirement that the magnetic field is parallel to the electric field. This selection of potentials leads to the following Hamiltonian
@@
where $E$ is the energy of the system. Then, one can use these operators to write down a couple of eigenvalue equations
@@
where $\delta z$ and $\delta y$ are constants. Both expressions will lead to valid solutions of the system. The simplest one is Eq.(50), it leads to a solution of the form
@@
${\hat{U}}_{z}=\exp\left(-i\frac{\delta y}{\hbar}{\hat{p}}_{z}\right),$ (61)
@@
V Conclusions
For both autonomous systems presented, it was shown that time-dependent conserved operators can be found and used to solve the Schrödinger equation. The degeneracy of the systems was analyzed, using the conserved operators as generators of solutions of the partial differential equation. Then, these operators were used to define the symmetries of the systems. It was proved that the invariance of the wave function under the action of the unitary operator, $\hat{U}_{x}$, leads to the quantization of resistance in integer multiples of Klitzing’s constant, in agreement with experimental results [4, 5, 6]. Finally, it is worth noting that Tao and Wu already pointed out that the degeneracy of the system is crucial for quantized resistance [20], a result that aligns with the findings presented in this work.
@@ Acknowledgment
This work was supported by DGAPA-UNAM Posdoctoral Program (POSDOC).

### v3

@@ Abstract
Two systems for a charged particle are studied using an adaptation of Lewis and Riesenfeld invariant method. The first system involves a particle under the effect of a constant electric field, and the second system adds a parallel magnetic field. In both cases, time-dependent conserved operators are identified, which can be used to derive time-dependent wave functions for the Schrödinger equation, where the time variable is not separable from the space coordinates. These conserved operators are also used to construct unitary operators, which define the symmetries of the systems. Finally, it is shown that the invariance of the wave function under the action of these unitary operators leads to the quantization of resistance equal to an integer number times the Klitzing’s constant.
Time-dependent operators, Resistance quantization, Klitzing’s constant, Constant force
@@ I Introduction
Conserved operators in dynamical systems have proven to be powerful tools in quantum mechanics for identifying suitable bases and understanding the structure of quantum states. A classical example is the central force problem, where the conservation of angular momentum allows one to determine the eigenfunctions associated with the azimuthal and polar angles [1]. Another well-known case involves a charged particle in a uniform magnetic field. By choosing Landau’s gauge, the linear momentum becomes a conserved quantity, leading to the so-called Landau ansatz. The corresponding eigenfunctions are products of plane waves and harmonic oscillator wave functions [2].
More recently, it has been shown that for a charged particle in a constant electromagnetic field, a complete description of the system requires not only Landau’s ansatz but also a time-dependent conserved operator [3]. In that work, it was demonstrated that requiring the wave function to be invariant under a unitary transformation generated by this conserved operator leads to the quantization of resistivity for a single particle. Interestingly, the resulting expression closely resembles the quantized resistance observed experimentally by Störmer [4, 5, 6]. Moreover, the invariance condition yields two distinct quantization rules: one for the magnetic flux (also derived through Landau’s ansatz [2]) and another associated with the electric field, expressed in Eq. (36). The latter serves as the primary motivation for the present study.
@@
In this work, we analyze two autonomous quantum systems that admit time-dependent conserved operators and naturally lead to the quantization condition for the electric field. The first system describes a particle subjected to a constant electric field, while the second considers a particle in a constant electromagnetic field with parallel electric and magnetic components. The linear potential involved in the first case is particularly relevant in studies of energy level shifts (Stark effect) [7, 8, 9, 10], and time-dependent solutions have previously been obtained using the generalized Baker-Campbell-Hausdorff formula for arbitrary initial conditions [11]. However, the approach adopted here is based on the conserved quantities of the system, employing the Lewis–Riesenfeld invariant method which has proven to be a powerful tool to address time-dependent Hamiltonians [12, 13, 14, 15, 16, 17, 18, 19]. Although originally developed for non-autonomous systems [20, 21, 22], we show that this method can also be effectively applied to time-independent (autonomous) Hamiltonians.
The second configuration has been explored in various contexts. Classically, it has been used to study photodetachment from negative ions and the spatial distribution of monoenergetic electrons [23, 24, 25]; semiclassically, it has played a key role in photoionization microscopy experiments [26, 27]; and in fully quantum treatments, it has been employed to investigate resonances in non-hydrogenic Rydberg atoms, hydrogen ground and excited states, lithium atoms, and atomic photoionization spectra [28, 29, 30, 31].
Beyond these specific applications, the central result of this work is the demonstration that time-dependent wave functions in the studied systems require the resistance to be quantized, provided they remain invariant under a unitary transformation generated by a time-dependent conserved operator. In particular, the resistance takes the form $R=hn/q^{2}$, where $n\in\mathbb{N}$, making it quantized in units of the von Klitzing constant $h/q^{2}$ [32, 33]. Notably, this expression mirrors the quantized resistance measured in the experiments by Störmer that led to the discovery of the Fractional Quantum Hall Effect (FQHE) [4]. However, in contrast to the FQHE—which arises from electron–electron interactions in a two-dimensional electron gas under strong magnetic fields—the quantization presented for the first system emerges from a purely single-particle system constrained to one dimension and influenced only by a constant electric field. This suggests that similar forms of resistance quantization can arise from fundamentally different physical mechanisms. In the second system considered, although it is two-dimensional, the Hamiltonian still describes a single-particle scenario, and the same quantized expression is recovered—indicating that the quantization persists along the direction in which the electric field is applied. Interestingly, recent studies have shown that single-particle spectral features can serve as experimental signatures of Fractional Quantum Anomalous Hall Effect (which is the observation of FQHE without a magnetic field) states in van der Waals heterostructures, highlighting the intriguing possibility that FQHE signatures may be accessible through single-particle measurements [34].
II invariant method
In this section, the invariant method adapted to autonomous systems with electromagnetic field is reviewed briefly. The systems addressed in this work are described by the non-relativistic spinless particle in an constant electromagnetic field such that the Hamiltonian, written in centimetre-gram-second (CGS) units, is
@@ III One-dimensional constant electric field
$\hat{p}_{x}\psi-q{\cal E}t\psi=\lambda\psi,$ (12)
where $\lambda$ are the eigenvalues of the operator. These eigenvalues can be interpreted as a variation of time, the above equation can be rewritten as
$\hat{p}_{x}\psi=q{\cal E}\left(t+\frac{\lambda}{q{\cal E}}\right)\psi,$ (13)
and defining the variation as $\delta t=\pm\lambda/q{\cal E}$ having second as units, in this particular case the negative sign was chose and the results do not depend on the selection of this sign. Then, the eigenvalue equation reads as follows
@@
Solving the above equation, it can be found that
@@
where the normalization constant was calculated over the finite interval $x\in[-L/2,L/2]$, which can be interpreted as placing the particle within a uniform electric field generated by a parallel-plate capacitor. Although the potential formally extends over the entire real line, this restriction provides a physically motivated and mathematically convenient framework that avoids divergences in the normalization process. At this point, one can search for the simplest form of the solution, in a sense that the above wave function is just a unitary transform of the new function. The conserved operators in Eq.(10) and Eq.(11) can be used to define a couple of unitary operators
@@
where $\delta x$ is a constant having units of length. These operators define the symmetries of the system such that the Schrödinger equation remains invariant under a unitary transformation, i.e.,
@@
The quantized resistance derived above exhibits noteworthy characteristics. It is directly proportional to an integer, reminiscent of the quantization observed in the FQHE, first measured by Tsui, Störmer, and Gossard [4]. It is well established that the FQHE arises from strong electron-electron interactions in a two-dimensional electron gas subjected to a strong magnetic field [36]. Nevertheless, exploring alternative mechanisms for realizing FQHE-like quantized resistance remains an active area of research [37, 38, 39, 40, 41, 42], particularly in graphene and multilayer graphene lattices [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. In contrast with the electron-electron interaction FQHE hypothesis, the quantization eq.(38) emerges from a single-particle model confined to one dimension, without invoking many-body effects or the presence of a magnetic field. This highlights the fundamentally different physical origin of the quantization observed here.
To finalize, an analogy with classical mechanics can be made. Using the wave function in eq. (25) and substituting it into the definition of the quantum current,
$J=\frac{i\hbar}{2m}\left(\varphi\frac{\partial\varphi^{*}}{\partial x}-\varphi^{*}\frac{\partial\varphi}{\partial x}\right),$ (39)
one can obtain
$J=\frac{q{\cal E}t}{m}|\varphi|^{2}.$ (40)
Since the current is, by definition, the velocity $v$ times the particle density $\rho=|\varphi|^{2}$, i.e., $J=v\rho$, from the above equation, the velocity is given by
$v=\frac{q{\cal E}t}{m},$ (41)
and, differentiating with respect to time and multiplying by the mass, one can recover the classical Newton force for a constant electric field:
$m\frac{dv}{dt}=q{\cal E}.$ (42)
@@ IV Constant electromagnetic field
As mentioned in the introduction, another autonomous system with a time-dependent conserved operator is the case of a constant electromagnetic field. The analysis for the situation where the electric field is perpendicular to the magnetic field can be found in reference [3]. Here, it will be consider the scenario where the electric field is parallel to the magnetic field. In this case, the same electric potential than in the previous section is used. On the other hand, to consider the magnetic field in Hamiltonian Eq.(1), it is necessary to define the vector potential. In this case ${\bf A}=B(0,0,y)$ satisfies the requirement that the magnetic field is parallel to the electric field. This selection of potentials leads to the following Hamiltonian
@@
where $E$ is the energy of the system. Then, one can use these operators to write down a couple of eigenvalue equations and similar to the previous case the eigenvalues have the effect to displace the solutions by a quantity denoted as $\delta y$ and $\delta z$, such that allow us to write down the equations as
@@
where $\delta z$ and $\delta y$ are constants having length as units. Both expressions will lead to valid solutions of the system. The simplest one is Eq.(56), it leads to a solution of the form
@@
${\hat{U}}_{z}=\exp\left(-i\frac{\delta z}{\hbar}{\hat{p}}_{z}\right),$ (67)
@@
Conclusions
In this work, it was demonstrated that time-dependent conserved operators can be used to solve the Schrödinger equation for two autonomous systems, applying the Lewis–Riesenfeld invariant method. The degeneracy of the systems was explored, revealing how these conserved operators act as generators of solutions to the corresponding partial differential equation and how they define the symmetries of the systems. The invariance of the wave function under the unitary transformation $\hat{U}_{x}$ leads to resistance quantization, expressed as integer multiples of the von Klitzing constant.
The condition $\hat{U}_{x}\Psi=\Psi$ suggests a scenario in which all electrons in a closed system initially occupy the same quantum state. After a displacement by $\delta x$, the wave function returns to its initial form, demonstrating that this symmetry is necessary for quantization.
Finally, the quantized resistance given by Eq. (38) closely resembles the form observed in the Störmer experiment. While the FQHE is commonly attributed to strong electron-electron interactions in two-dimensional systems under high magnetic fields, the results presented here show that similar resistance quantization can arise in much simpler, non-interacting single-particle models. This emphasizes that symmetry and conserved quantities alone can lead to quantized behavior reminiscent of the FQHE—even in the absence of interactions and, in the first system, under the influence of an electric field only.
@@ Acknowledgment
This work was supported by DGAPA-UNAM Posdoctoral Program (POSDOC), the project UNAM-PAPIIT IG100725 and project SECIHIT Fronteras CF-2023-G. I am grateful for these supports.